# Review of Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Industrial Open Source Data

### v1

@@ 1 Introduction
In the contemporary landscape of Industry 4.0, the emphasis on the reliability, efficiency, and longevity of industrial systems has become crucial [34]. Against this backdrop, Prognostics and Health Management (PHM) has crystallized as an indispensable discipline, that integrates the detection, diagnosis, assessment, and prognosis of system failures addressing the growing need for proactive system health management [82]. This integrative approach can bolster the reliability and safety of industrial engineering systems, with minimal unplanned downtimes and reduced maintenance costs.
@@
The rise of the Internet of Things [62], big data analytics [74], cyber-physical systems [35], machine learning (ML) [21], deep learning (DL) [57], and industrial artificial intelligence [54, 36] has paved the way for a transformative shift in PHM. Historically, PHM strategies largely relied on physical-based methods, which derived their strength from profound insights into system physics, material properties, and failure mechanisms. However, these strategies often struggled with scalability, adaptability, and handling the vast variability and uncertainties inherent in real-world operations. Data-driven approaches, in contrast, underpinned by machine learning algorithms, offer the ability to model complex systems, identify intricate patterns, and diagnose and predict failures with high accuracy. This evolution underscores the potential of data-driven models to deliver precise and adaptable diagnostic and prognostic solutions, even in the absence of explicit physical models.
@@
Over the last decade, organizations such as the PHM Society and the IEEE Reliability Society have played pivotal roles in fostering innovation, research, and collaboration in the PHM field. One notable initiative is the PHM data challenge competition, which stands as a testament to the growing emphasis on real-world, data-centric PHM solutions. By providing participants with genuine industrial datasets and posing real-world challenges, this competition seeks to accelerate the development and validation of cutting-edge PHM methodologies, bridging the gap between academia and industry. In light of the recent surge in data-driven methodologies, particularly machine learning, and deep learning, in PHM, it becomes essential to revisit and scrutinize research algorithms and methodologies derived from the PHM data challenge competitions and their datasets. Systematically reviewing the challenges presented by industrial data across various sectors with methodologies and analytics can foster and enhance a deeper understanding of data-driven approaches. Such endeavors not only propel the evolution of data-centric techniques, machine learning, and artificial intelligence (AI) within the PHM domain but also accentuate advancements in the broader landscape of industrial data analytics and industrial artificial intelligence.
@@
This study aims to conduct a comprehensive problem-challenge-solution-oriented review within the domain of PHM Data Challenge. Specifically, this study summarizes the problems and challenges presented in nine PHM data challenge competitions, elucidating the algorithms, methods, and analytical techniques employed to address them. Further, this study encapsulates the common challenges associated with industrial data, underscoring specific issues related to data issues, and modeling in data-driven approaches. In pursuit of the objective, approximately sixty research papers were reviewed, encompassing both competition-winning contributions and subsequent exploratory research undertaken post-competition.
The rest of this paper is organized as follows: Section 2 provides an overview of PHM data challenge competitions and underscores the principal research tasks within PHM. Section 3 and Section 4 respectively introduce the prevalent challenges associated with two pivotal tasks: detection & diagnosis, and assessment & prognosis, subsequently detailing the solutions presented in various PHM data competitions individually. Section 5 critically summarizes and examines common challenges in the PHM domain through the lens of data-driven methodologies, addressing concerns related to data issues, and model issues. Section 6 provides a forward-looking perspective on the enhanced integration of data-driven techniques, machine learning, and AI in future PHM endeavors. Section 7 concludes the paper, highlighting its findings and contributions.
2 Background
2.1 Overview of PHM Data Challenge Competitions
From 2018 to 2023, PHM Society and IEEE Reliability Society have initiated nine PHM data challenge competitions which reflect challenges associated with analyzing industrial data across diverse industrial sectors. The main purpose behind these competitions is to amplify focus and collaborative efforts from both academic and industrial realms to address real-world challenges. The topics of these competitions are commendably extensive, encompassing diverse industrial entities such as Ion Mill Etching Tools, Filtration Systems, Manufacturing Production Lines, Turbofan Engines, Printed Circuit Boards, Rock Drills, Spacecraft Propulsion Systems, and Gearbox, among others. Furthermore, the range of problems posed encapsulates the main tasks in the PHM domain including detection, diagnosis, assessment, and prognosis. These tasks align with PHM’s ultimate goal: to accurately forecast system health, degradation, and eventual failure, thereby improving system reliability, safety, and operational efficiency. For ease of discussing different competitions, we’ve adopted a condensed naming convention for the PHM data challenge competition: ”YEAR ORGANIZATION” indicates the respective year and organizing body of the challenge. Among the challenges discussed, one is organized by IEEE Reliability Society, while the remaining eight are conducted by PHM Society. It’s noteworthy that the PHM Society conducts an annual conference in North America, an Asia-Pacific conference in odd years, and a European conference in even years. We use the abbreviations ”PHM NA”, ”PHM AP”, and ”PHM EU” to represent the competitions held in North America, Asia-Pacific, and Europe, respectively. For instance, ”2018 PHM NA” refers to the challenge held in North America by the PHM Society in 2018 while ”2023 IEEE” represents the competition hosted by the IEEE Reliability Society in 2023. Table 1 presents the systems, research tasks, and common challenges of the nine data competitions. For an in-depth overview of the PHM data challenge competitions and their associated datasets, please refer to the appendix.
@@ 2.1 Overview of PHM Data Challenge Competitions
Table 1: Overview of PHM Data Challenge Competitions from 2018 to 2023
@@ 2.2 Core Research Tasks within PHM
Assessment: Following the identification of the root cause, the assessment phase evaluates the current operational status and performance of the system. Leveraging either historical data or recent machinery behavior, this phase evaluates potential risks or assesses the health status of the system in its present condition.
Prognosis: Representing the most forward-looking phase, prognosis leverages both current and historical data to forecast the future health of a system or the residual life of a machine. Commonly, this is referred to as predicting the Remaining Useful Life (RUL). This pivotal phase provides estimates on potential system or machinery failure timelines, thereby facilitating proactive maintenance planning. To achieve this, data-driven approaches, including machine learning and artificial intelligence, are often employed to construct predictive models for future prognosis.
Figure 1: Research Tasks in PHM Data Challenge Competitions.
@@ 2.3 Comprehensive Review of Data-Driven Approaches in Recent PHM Data Challenge Competitions
Figure 2: A Unified Machine Learning Framework in Prognostics and Health Management Domain.
@@
After conducting an in-depth review of data competitions from the last six years, we have summarized a unified machine learning framework that concludes machine learning approaches for PHM data competitions during this period. As shown in Figure 2, it encompasses five primary components: Data Collection, Data Processing, Data Visualization, Conventional Machine Learning & Deep Learning, and Model Interpretability.
@@
Data Collection: Data collection is the foundation of these open-source data challenge competitions. Companies across various industries contribute datasets with the objective of harnessing data-driven strategies to address intricate challenges in PHM. Typically, these industrial datasets encompass a diverse range of data types, including but not limited to time-series data, tabular data, images, sensor readings, and simulation data.
@@
Data Processing: A pivotal stage before ML model development is data processing, crucial for enhancing data quality and ensuring effective model training. This stage can be further subdivided into two aspects: data preprocessing and feature engineering. Data preprocessing addresses raw data challenges, typically including missing data, noise, outliers, data imbalance, and scaling issues. Techniques like imputation, denoising, outlier detection, resampling, and normalization/standardization are commonly applied. Feature engineering follows, refining data representation post-preprocessing. This step often uncovers hidden insights and deepens understanding of data, therefore significantly enhancing model performance and predictive capabilities in the later stage.
@@
Data Visualization: Data visualization is another important aspect of the machine learning process. It involves transforming data into intuitive graphical representations, such as graphs or charts. Researchers can use these charts to better observe trends, patterns, or outliers in data, which further help people have a better understanding of data and generate some useful insights. In the PHM domain, effective visualization can help initial data exploration and analysis, and accelerate data processing and ML modeling development.
@@
Conventional Machine Learning & Deep Learning Techniques: Post data processing, various conventional machine learning, and deep learning methods are developed to solve the competition problems using useful and usable data, as shown in Figure 2. This segment includes ”model training and testing” and ”model prediction and classification”. During the training and testing phase. Observed techniques in these competitions are diverse, including Tree-Based Methods, CNNs, RNNs and their variants, SVMs, Unsupervised Learning, Ensemble Learning, Similarity-Based Methods, Rule-Based Methods, Transfer Learning, Domain Adaptation, etc. Once the machine learning models are trained and tested well, they become invaluable tools for various PHM applications which include detecting abnormal performance or faults, classifying different root cause faults, assessing the current health state of systems, or predicting the remaining useful life of components. These applications are not only critical in meeting the diverse requirements of PHM data challenge competitions but also demonstrate the adaptability and effectiveness of the various machine learning and deep learning approaches utilized.
@@
Model Interpretability: Except for the pursuit of accuracy in predictions and classifications, model interpretability is becoming more and more important. Model interpretability refers to interpreting machine learning model outputs [52] which helps humans understand the ’why’ behind a model’s predictions, facilitating collaboration and more informed decision-making. Many explainable artificial intelligence (XAI) methods such as Local Interpretable Model-Agnostic Explanations (LIME) [58] and SHapley Additive exPlanations (SHAP) [48] are utilized in the PHM data challenge competitions to explain model outputs in PHM [66].

### v2

@@ 1 Introduction
In the era of Industry 4.0, the emphasis on the reliability, efficiency, and longevity of industrial systems has become crucial [53]. Against this backdrop, Prognostics and Health Management (PHM) has emerged as an indispensable discipline, that integrates the detection, diagnosis, assessment, and prognosis of system failures addressing the growing need for proactive system health management [127]. This integrative approach can bolster the reliability and safety of industrial systems, with minimal unplanned downtimes and reduced maintenance costs.
@@
The rise of the Internet of Things [94], big data analytics [110], cyber-physical systems [55], machine learning (ML) [37], deep learning (DL)  [87], and industrial artificial intelligence [81, 57] has paved the way for a transformative shift in PHM. Historically, PHM strategies largely relied on physical-based methods, which derived their strength from profound insights into system physics, material properties, and failure mechanisms. These approaches, however, often struggled with scalability, adaptability, and the ability to handle the vast variability and uncertainties inherent in real-world operations. In contrast, ML techniques present the capability to model complex systems comprehensively, uncover intricate patterns, and accurately diagnose and predict failures. Yet, employing ML in PHM poses several challenges, including the critical need for usable, useful, high-quality, and extensive data sets, extensive computing resources, a solid infrastructure for data collection and processing, as well as the necessity for assembly of a team of professionals proficient in artificial intelligence (AI) and ML [56, 5]. Despite these challenges, the momentum for implementing ML in industrial settings is evident. With the gradual move towards digital transformation within the industry, incorporating ML into PHM aligns with the principles of Industry 4.0 [82]. It underscores the capability of data-driven approaches to provide precise and adaptable diagnostic and prognostic solutions and represents a strategic shift towards leveraging ML techniques for enhanced predictive maintenance.
1.1 A Survey of Relevant Reviews
A multitude of comprehensive reviews on ML strategies for PHM are noted [90, 83, 52, 79, 122, 78, 123, 82, 87]. These reviews evaluate the existing literature both qualitatively and quantitatively, presenting their distinct perspectives and pinpointing the trends and new concepts of ML strategies for PHM across various scenarios. [90] reviewed DL architectures like a one-class neural network (OCNN), self-organizing map (SOM), and generative techniques aligned with industrial needs from a predictive maintenance perspective. [83] highlighted seven DL architectures, including emerging DL methods such as graph neural networks, transformers, and generative adversarial networks, addressing four different challenges (imbalanced data, multimodal data fusion, compound fault types, and edge device implementation). Additionally, [52] provided a comprehensive examination of PHM strategies in the context of smart factories, spanning from traditional approaches to DL-based approaches. An extensive review on modeling techniques supporting PHM of industrial equipment, specifically within onshore wind energy and civil aviation sectors, was given by [122], wherein they discuss how modeling approaches are shaped by industry-specific factors (maintenance strategies, implementation aspects, and supporting technologies). Furthermore, [78] proposed a general guideline under AI-based PHM for selecting appropriate techniques to solve specific PHM problems.
The aforementioned review papers serve as the groundwork for further PHM development. Currently, a significant portion of the literature primarily examines algorithms based on their capabilities and functionalities, with numerous reviews covering various ML or DL methods like CNN, RNN, GAN, GNN, Transformer, etc. However, how these algorithms are applied within actual industrial settings is still rarely considered. The practice among many researchers of relying on artificial data sets for algorithm testing, instead of using real, industry-specific data sets, is notable. Furthermore, a considerable portion of the data sets mentioned in these studies are not publicly accessible, and among the available ones, most date back more than six years (before 2018). Additionally, the focus of discussed data sets predominantly lies on mechanical components commonly used in algorithm development, such as gears and bearings. In light of these observations, this paper will try to analyze the open source data sets available in the last six years and conduct an in-depth analysis from the perspective of problem-challenge-solution-application orientation.
1.2 Motivation
@@
Over the last fifteen years, organizations such as the PHM Society and the IEEE Reliability Society have played pivotal roles in fostering innovation, research, and collaboration in the PHM field. One notable initiative is the PHM data challenge competition, which stands as a testament to the growing emphasis on real-world, data-centric PHM solutions. By providing participants with real industrial datasets and posing real-world challenges, this competition seeks to accelerate the development and validation of cutting-edge PHM methodologies, bridging the gap between academia and industry. In light of the recent surge in data-driven methodologies, particularly ML, and DL, in PHM, it becomes essential to revisit and scrutinize research algorithms and methodologies derived from the PHM data challenge competitions and their datasets. Systematically reviewing the challenges presented by industrial data across various sectors with methodologies and analytics can foster and enhance a deeper understanding of data-driven approaches. Such endeavors not only propel the evolution of data-centric techniques, ML, and AI within the PHM domain but also accentuate advancements in the broader landscape of industrial data analytics and industrial artificial intelligence.
After evaluating competitions over the past 15 years, we chose to concentrate on the competitions held after 2017, specifically from 2018 to 2023, covering a span of six years. This decision is motivated by several factors. Firstly, as we mentioned in section 1.1, the open-source datasets that are utilized for analytics in the aforementioned review papers are most before 2018, while the post-2018 open-source datasets are rarely discussed in those reviews. Secondly, the PHM data challenge competition offers open-source datasets, facilitating deeper investigation by researchers into the challenges prevalent across various industrial settings. This, in turn, encourages the development of innovative algorithms and fosters advancements in the PHM field. Thirdly, recent years have witnessed rapid advancements and iterations in ML and DL technologies, leading to the emergence of a large number of complex architectures and algorithms. Reviewing these papers can help to better understand the latest trends of ML and DL in PHM and their role in the future development of PHM.
1.3 Contributions
@@
This study aims to conduct a comprehensive problem-challenge-solution-application-oriented review within the domain of PHM Data Challenge. In pursuit of the objective, approximately sixty research papers were reviewed, encompassing both competition winning contributions and subsequent exploratory research undertaken post-competition. Our contributions include the following:
Table 1: Overview of PHM Data Challenge Competitions from 2018 to 2023
@@
  1. 1.
    This study summarizes the problems and solutions presented in nine PHM data challenge competitions, elucidating the ML or DL algorithms, methods, and analytical strategies employed to tackle these competitions.
  2. 2.
    We propose a unified ML framework for the PHM domain based on this review study, serving as a general guideline for the development of future ML models.
  3. 3.
    We discuss common challenges associated with industrial data, underscoring specific issues related to data issues (missing data, data imbalance, and domain shift), and model issues (model selection, machine learning model interpretability, model robustness and generalization) in data-driven approaches. Possible solutions are also provided. Moreover, we evaluate the limitations of these competitions and suggest future directions for their evolution.
  4. 4.
    We identify six further research directions for the application of ML in PHM. These include (1) a need for open-source multi-modal datasets, (2) development of multi-modal machine learning approaches, (3) further exploration in machine learning model interpretability, (4) novel transfer learning and domain adaptation techniques development for model robustness and generalization, (5) construction of large knowledge library, and (6) potential utilization of large language models and industrial large knowledge models.
The rest of this paper is organized as follows: Section 2 provides an overview of PHM data challenge competitions, a unified ML framework for the PHM domain, and underscores the principal research tasks within PHM. Section 3 and Section 4 respectively introduce the prevalent challenges associated with two pivotal tasks: detection & diagnosis, and assessment & prognosis, subsequently detailing the solutions presented in various PHM data competitions individually. Section 5 critically summarizes and examines common challenges in the PHM domain through the perspective of ML methodologies, addressing concerns related to data issues, and model issues. The limitations of these competitions are also examined. Section 6 provides a forward-looking perspective on the enhanced integration of data-driven techniques, ML, DL, and AI in future PHM development. Section 7 concludes the paper, highlighting its findings and contributions.
2 Background
2.1 Overview of PHM Data Challenge Competitions
From 2018 to 2023, PHM Society and IEEE Reliability Society have initiated nine PHM data challenge competitions which reflect challenges associated with analyzing industrial data across diverse industrial sectors. The main purpose behind these competitions is to amplify focus and collaborative efforts from both academic and industrial realms to address real-world challenges. The topics of these competitions are commendably extensive, encompassing diverse industrial systems such as Ion Mill Etching Tools, Filtration Systems, Manufacturing Production Lines, Turbofan Engines, Printed Circuit Boards, Rock Drills, Spacecraft Propulsion Systems, and Gearbox, among others. Furthermore, the range of problems posed encapsulates the main tasks in PHM domain including detection, diagnosis, assessment, and prognosis. These tasks align with PHM’s ultimate goal: to accurately forecast system health, degradation, and eventual failure, thereby improving system reliability, safety, and operational efficiency.
For ease of discussing different competitions, we’ve adopted a condensed naming convention for the PHM data challenge competition: ”YEAR ORGANIZATION” indicates the respective year and organizing body of the challenge. Among the challenges discussed, one is organized by the IEEE Reliability Society, while the remaining eight are organized by the PHM Society. The PHM Society conducts an annual conference in North America, an Asia-Pacific conference in odd years, and a European conference in even years. We use the abbreviations ”PHM NA”, ”PHM AP”, and ”PHM EU” to represent the competitions held in North America, Asia-Pacific, and Europe, respectively. For instance, ”2018 PHM NA” refers to the challenge held in North America by the PHM Society in 2018 while ”2023 IEEE” represents the competition hosted by the IEEE Reliability Society in 2023. Table 1 presents the systems, research tasks, and common challenges of the nine data competitions. For an in-depth overview of the PHM data challenge competitions and their associated datasets, please refer to Appendix, Section 3 and Section 4.
@@ 2.2 Core Research Tasks within PHM
Assessment: In the assessment phase, the current operational status and performance of the system will be evaluated. Leveraging either historical data or recent machinery behavior, this phase evaluates potential risks or assesses the health status of the system in its present condition.
Prognosis: Representing the most forward-looking phase, prognosis leverages both current and historical data to forecast the future health of a system or the residual life of a machine. Commonly, this is referred to as predicting the Remaining Useful Life (RUL). This pivotal phase provides estimates on potential system or machinery failure timelines, thereby facilitating proactive maintenance planning. To achieve this, data-driven approaches, including ML and AI, are often employed to construct predictive models for future prognosis.
@@
Figure 1: Research Tasks in PHM Data Challenge Competitions
@@
Figure 2: Distribution of ML Approaches Across PHM Data Challenge Competitions
@@ 2.3 Comprehensive Review of Data-Driven Approaches in Recent PHM Data Challenge Competitions
After conducting an in-depth review of data competitions from the last six years, Figure 3 illustrates the publication count and the density of specific ML or DL approaches discussed in this paper. Furthermore, Figure 2 provides details into the frequency of particular ML approaches within each competition. It is important to note that we count the occurrences of distinct ML methods mentioned in a research paper. Given that some articles employ multiple ML methods, the aggregate count of methods exceeds the total number of publications.
Figure 3: The Density of Various ML Approaches in PHM Data Challenge Competitions from 2018 until December 2023
@@
Moreover, we have summarized a unified ML framework that concludes ML approaches for PHM data competitions during this period. As shown in Figure 4, it encompasses five primary components: Data Collection, Data Processing, Data Visualization, Conventional Machine Learning & Deep Learning, and Model Interpretability.
Figure 4: A Unified Machine Learning Framework in Prognostics and Health Management Domain
@@
Data Collection: Data collection is the foundation of these open-source data challenge competitions (PHM Society [^1] https://phmsociety.org/ and IEEE Reliability Society [^2] https://rs.ieee.org/). Companies across various industries contribute datasets with the objective of harnessing data-driven strategies to address intricate challenges in PHM. Typically, these industrial datasets encompass a diverse range of data types, including but not limited to time-series data, tabular data, images, sensor readings, and simulation data.
@@
Data Processing: A pivotal stage before ML model development is data processing, which is crucial for enhancing data quality and ensuring effective model training [17, 105]. This stage can be further subdivided into two aspects: data preprocessing and feature engineering. Data preprocessing addresses raw data challenges, typically including missing data, noise, outliers, data imbalance, and scaling issues. Techniques like imputation [28], denoising [22], outlier detection [72], resampling [16], and normalization & standardization [54, 33] are commonly applied. Feature engineering follows, refining data representation post-preprocessing. This step often uncovers hidden insights and deepens understanding of data, therefore significantly enhancing model performance and predictive capabilities in the later stage [92].
@@
Data Visualization: Data visualization is another important aspect of the ML process. It involves transforming data into intuitive graphical representations, such as graphs or charts. Researchers can use these charts to better observe trends, patterns, or outliers in data, which further help people have a better understanding of data and generate some useful insights. In the PHM domain, effective visualization can help initial data exploration and analysis, and accelerate data processing and ML modeling development [12, 15].
@@
Conventional Machine Learning & Deep Learning Techniques: Post data processing, various conventional ML, and DL methods are developed to solve the competition problems using useful and usable data, as shown in Figure 4. This segment includes ”model training and testing” and ”model prediction and classification”. During the training and testing phase. Observed techniques in these competitions are diverse, including Tree-Based Methods [22, 102, 43], CNNs [69, 24, 62], RNNs and their variants [7, 62, 48, 107], Transformers [111, 124], DANN [49], SVMs [75, 80], Unsupervised Learning, Ensemble Learning [86, 107], Similarity-Based Methods [73, 47, 2], Rule-Based Methods [3, 61], Transfer Learning [66, 120, 8], Domain Adaptation [80, 49], etc. Once the ML models are trained and tested well, they become invaluable tools for various PHM applications which include detecting abnormal performance or faults, classifying different faults, assessing the current health state of systems, or predicting the RUL of components. These applications are not only critical in meeting the diverse requirements of PHM data challenge competitions but also demonstrate the adaptability and effectiveness of the various ML and DL approaches utilized.
@@
Model Interpretability: Except for the pursuit of accuracy in prediction and classification problems, model interpretability is becoming more and more important. Model interpretability refers to interpreting ML model outputs [76] which helps humans understand the ’why’ behind a model’s predictions, facilitating collaboration and more informed decision making. Many explainable artificial intelligence (XAI) methods such as Local Interpretable Model-Agnostic Explanations (LIME) [88] and SHapley Additive exPlanations (SHAP) [70] are utilized in the PHM data challenge competitions to explain model outputs in PHM [98].
@@ 3 Detection & Diagnosis
Table 2: Overview of Detection and Diagnosis Problems in PHM Data Challenge Competitions
2021 PHM EU 2022 PHM EU 2022 PHM NA 2023 PHM AP 2023 IEEE
Failure Mode 8 1+1+2 10 3 4
Sensor Number 50 NA 3 7 1
Asset NA NA 8 4 NA
Operating Condition 2 1+1+1 1 1 2
Volume of Data Limited Medium Large Limited Large
(70 Train, 29 Test) (SPI:1924, AOI:1924) (37229 Train, 16396 Test) (177 Train, 46 Test) (Total 50000)
Sampling Rate 0.1 Hz NA 50 kHz 1 kHz 10 kHz
Sampling Interval 1-3 hours NA NA 1.2 s 5 minutes